OpenStudy (anonymous):

Find the next three terms of the sequence. Then write a rule for the sequence. 648, 216, 72, 24

OpenStudy (anonymous):

my answer was 8,8/3,8/9 x(n)=x(n-1)/3

OpenStudy (anonymous):

yes, that looks right good job :)

OpenStudy (anonymous):

youre right with the sequence however rule will be a/(3^(n-1)) where a = 648

OpenStudy (anonymous):

and nice recursive formula

OpenStudy (anonymous):

but my teacher says its wrong or incomplete

OpenStudy (anonymous):

maybe your teacher wants you to simplify 8/3? maybe your teacher wants you to give the formula @Meinme gave

OpenStudy (anonymous):

@LilySwan as i said the formula u dervied wont work everywhere in the series

OpenStudy (anonymous):

its a geometric progression

OpenStudy (anonymous):

yeah its not Arithmetic progression its geometric because theres a common ratio not common difference

OpenStudy (anonymous):

so @LilySwan do you know how to find the equation of the sequence?

OpenStudy (anonymous):

my teacher commented this You've given me a recursive formula. For your rule think about what you are doing each time. How should that be expressed in the rule?

OpenStudy (anonymous):

yes, that makes sense. what the teacher wants you to do is find the formula that applies your recursive formula to the first value a variable number of times

OpenStudy (anonymous):

right!!

OpenStudy (anonymous):

the general formula of a geometric sequence is a*r^(n-1)

OpenStudy (anonymous):

a is the first term r is what you multiply by each time to get the next term, this is what you found in your recursive formula n is which term it is, (it is n-1 because a is the first term instead of term 0)

OpenStudy (anonymous):

are you confused?

OpenStudy (anonymous):

very much

OpenStudy (anonymous):

sorry

Directrix (directrix):

@LilySwan Is r = 1/3 ? Is the common ratio 1/3? Each term is 3 times the one preceeding it? So, 216 = 1/3*648 and so on?

OpenStudy (anonymous):

this should help you be less confused

OpenStudy (anonymous):

@Directrix yes, we already got that

OpenStudy (anonymous):

@LilySwan does the link help?

OpenStudy (anonymous):

a bit

Directrix (directrix):

g(n) = 648* (1/3)^(n-1) = 2^3 * 3^4 * [(3)^(-1)]^(n-1) = 2^3*3^4*((3^(-1))^(n-1)= 2^3 * 3^(4) * ((3^(-1(n-1)) = 2^3 * 3 ^ (4 + (-n+1) = 8 * 3 ^ (5 - n) @LilySwan 8 * 3 ^ (5 - n) --> This is probably what the teacher wants to see.-->